An electric charge q caused an electric potential of V = -218 V at point A 10.6 m from q. What is the electric charge q?


1
Expert's answer
2022-05-22T14:37:39-0400
"V=k\frac{q}{r}"

"q=Vr/k\\
=-218*10.6/9*10^9=-2.57*10^{-7}\:\rm C"